### v5

# A Simple and Effective Reinforcement Learning Method for Text-to-Image Diffusion Fine-tuning

Abstract
[…]
1 Introduction
Diffusion models have emerged as a powerful tool for generative modeling (sohl2015deep; ho2020denoising), with a strong capacity to model complex data distributions from various modalities, like images (rombach2022high), text (austin2021structured), natural molecules (xu2023geometric), and videos (blattmann2023align).
Diffusion models are typically pre-trained on a large-scale dataset, such that they can subsequently generate samples from the same data distribution. The training objective typically involves maximizing the data distribution likelihood. This pre-training stage helps generate high-quality samples from the model. However, some applications might require optimizing a custom reward function, for example, optimizing for generating aesthetically pleasing images (xu2024imagereward), semantic alignment of image-text pairs based on human feedback (schuhmann2022laion), or generating molecules with specific properties (wang2024fine).
To optimize for such black-box objectives, RL-based fine-tuning has been successfully applied to diffusion models (fan2024reinforcement; black2023training; wallace2024diffusion; li2024aligning; gu2024diffusion). For RL-based fine-tuning, the reverse diffusion process is treated as a Markov decision process (MDP), wherein prompts are treated as part of the input state, the generated image at each time-step is mapped to an action, which receives a reward from a fixed reward model (environment in standard MDP), and finally the diffusion model is treated as a policy, which we optimize to maximize rewards. For optimization, typically PPO is applied (fan2024reinforcement; black2023training). In applications where getting a reward model is infeasible or undesirable, “RL-free” fine-tuning (typically offline) can also be applied (wallace2024diffusion). For this work, we only focus on diffusion model fine-tuning using “online” RL methods, specifically PPO (schulman2017proximal).
An advantage of PPO is that it removes the incentive for the new policy to deviate too much from the previous reference policy, via importance sampling and clipping operation (schulman2017proximal). While effective, PPO can have significant computational overhead. In practice, RL fine-tuning for diffusion models via PPO requires concurrently loading three models in memory: (i) The reference policy: The base policy, which is usually initialized with the pre-trained diffusion model. (ii) The current policy: The policy that is RL fine-tuned, and also initialized with the pre-trained diffusion model. (iii) The reward model: Typically, a large vision-language model, trained via supervised fine-tuning objective (lee2023aligning), which assigns a scalar reward to the final generated image during the online optimization stage. This can result in a considerable computational burden, given that each policy can potentially have millions of parameters. In addition to computational overhead, PPO is also known to be sensitive to hyper-parameters (engstrom2019implementation; zheng2023delve; huang2024n+).
Simpler approaches, like REINFORCE (williams1992simple) avoid such complexities, and could theoretically be more efficient. However, in practice, they often suffer from high variance and instability. A variant of REINFORCE: reinforce leave-one-out (RLOO(kool2019buy) has been proposed that samples multiple sequences per input prompt, and a baseline correction term to reduce the variance; however, it still suffers from sample inefficiency.
This raises a fundamental question about the efficiency-effectiveness trade-off in RL-based diffusion fine-tuning. In this work, first we systematically explore this trade-off between efficiency – a lower computational cost, and reduced implementation complexity (i.e., fewer hyper-parameters) – and effectiveness – stable training, and final performance. We compare a simple REINFORCE approach with the standard PPO framework, demonstrating that while REINFORCE greatly reduces computational complexity, it comes at the cost of reduced performance.
Motivated by this finding, we propose a novel RL for diffusion fine-tuning method, LOOP, which combines the best of the both worlds. To reduce the variance during policy optimization, LOOP leverages multiple actions (diffusion trajectories) and a (REINFORCE) baseline correction term per input prompt. To maintain the stability and robustness of PPO, LOOP leverages clipping and importance sampling.
Our approach is conceptually similar to the recently proposed GRPO method for RL fine-tuning of LLMs (shao2024deepseekmath). The key technical differences are: (i) LOOP does not apply standard-deviation normalization in the advantage calculation. Recent work on LLM fine-tuning suggests that removing this normalization term can improve performance (liu2025understanding). (ii) Following this recent work, LOOP omits the KL penalty term. Prior studies indicate that explicit KL regularization has minimal practical effect on performance (Black2023), and recent theoretical work shows that on-policy RL methods implicitly maintain KL proximity to the base policy even without explicit regularization (shenfeld2025rl). (iii) In the diffusion setting, the reverse process has a fixed sequence length across all generations, making sequence-length normalization unnecessary.
For the primary evaluation benchmark, we choose the text-to-image compositionality benchmark (T2I-CompBench; huang2023t2i. Text-to-image models often fail to satisfy an essential reasoning ability of attribute binding, i.e., the generated image often fails to bind certain attributes specified in the instruction prompt (huang2023t2i; ramesh2022hierarchical; fu2024enhancing). As illustrated in Figure 1, LOOP outperforms previous diffusion methods on attribute binding. As attribute binding is a key skill necessary for real-world applications, we choose the T2I-CompBench benchmark alongside two other common tasks: aesthetic image generation and image-text semantic alignment.
SD v2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DDPO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LOOP Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“White cat playing with black ball”
“Mechanical owl with cobalt blue feathers on a rusted bronze lamppost” “A hexagonal watermelon placed besides a knife”
“Black horse with glowing cyan patterns in maze of floating golden rings” “Neon pink cactus growing on a cobalt blue rock”
Figure 1: LOOP improves attribute binding. Qualitative examples presented from images generated via Stable Diffusion (SD) 2.0 (first row), DDPO (black2023training) (second row), and LOOP $k=4$ (third row). In the first prompt, SD and DDPO both fail to bind the color black with the ball in the image, whereas LOOP binds the color black to the ball. In the second example, SD and DDPO fail to generate rusted bronze color lamppost, whereas LOOP manages to do that. In the third image, SD and DDPO fail to bind the shape hexagon to the watermelon, whereas LOOP manages so. In the fourth example, SD and DDPO fail to generate the black horse with flowing cyan patterns, whereas LOOP generates the horse with the correct color attribute. Finally, in the last image, SD and DDPO fail to bind cobalt blue color to the rock, whereas LOOP binds that successfully.
To summarize, our main contributions are as follows:
  • PPO vs. REINFORCE efficiency-effectiveness trade-off. We systematically study how design elements like clipping, reference policy, value function in PPO compare to a simple REINFORCE method, highlighting the efficiency-effectiveness trade-off in diffusion fine-tuning. To the best of our knowledge, we are the first ones to present such a systematic study, highlighting the trade-offs in diffusion fine-tuning.
  • Introducing LOOP. We propose LOOP, a novel RL for diffusion fine-tuning method combining the best of REINFORCE and PPO. LOOP leverages multiple diffusion trajectories and a REINFORCE baseline correction term for variance reduction, as well as clipping and importance sampling from PPO for robustness and sample efficiency.
  • Empirical validation. To validate our claims empirically, we conduct experiments on the T2I-CompBench benchmark image compositionality benchmark. The benchmark evaluates the attribute binding capabilities of the text-to-image generative models and shows that LOOP succeeds where previous text-to-image generative models often fail. We also evaluate LOOP on two common objectives from the literature on RL for diffusion: image aesthetic and text-image semantic alignment (black2023training).
The remainder of the paper is organized as follows. In the next section, we provide the necessary background and discuss related work. Section 3 revisits the efficiency–effectiveness trade-off between REINFORCE and PPO. Section 4 introduces our proposed method, Leave-One-Out PPO (LOOP) for diffusion fine-tuning. Section 5 describes the experimental setup, and Section 6 presents the results and discussion. Finally, Section 7 concludes the paper.
2 Background and Related Work
2.1 Diffusion Models
We focus on denoising diffusion probabilistic models (DDPM) as the base model for text-to-image generative modeling (ho2020denoising; sohl2015deep). Briefly, given a conditioning context variable $\mathbf{c}$ (a text prompt in our case), and the data sample $\mathbf{x}_{0}$, DDPM models $p(\mathbf{x}_{0}\mid\mathbf{c})$ via a Markov chain of length $T$, with the following dynamics:
$p_{\theta}(\mathbf{x}_{0:T}\mid\mathbf{c})=p(\mathbf{x}_{T}\mid\mathbf{c})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{c}).$ (1)
Image generation in a diffusion model is achieved via the following ancestral sampling scheme, which is a reverse diffusion process:
$\mathbf{x}_{T}{}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\;\;\mathbf{x}_{t}\sim N\left(\mathbf{x}_{t}|\mu_{\theta}(\mathbf{x}_{t},\mathbf{c},t),\sigma^{2}_{\theta}I\right),\forall t\in[0,T-1],$ (2)
where the distribution at time-step $t$ is assumed to be a multivariate normal distribution with the predicted mean $\mu_{\theta}(\mathbf{x}_{t},\mathbf{c},t)$, and a constant variance.
2.2 Proximal Policy Optimization (PPO) for RL
[…]
PPO was introduced for optimizing a policy with the objective of maximizing the overall reward in the RL paradigm. PPO removes the incentive for the current policy $\pi_{t}$ to diverge from the previous policy $\pi_{t-1}$ outside the range $1-\epsilon,1+\epsilon]$, where $\epsilon$ is a hyper-parameter. As long as the subsequent policies are closer to each other in the action space, the monotonic policy improvement bound guarantees a monotonic improvement in the policy’s performance as the optimization progresses. This property justifies the clipping term in the mathematical formulation of the PPO objective function (schulman2015trust; achiam2017constrained; queeney2021generalized). Formally, the PPO objective function is:
$\mbox{}\hskip-5.69054ptJ(\theta)\!=\!\mathbb{E}\!\!\left[\min\!\left(\!r_{t}(\theta)\hat{A}_{t},\,\!\text{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\!\hat{A}_{t}\!\right)\!\right]\!,\!\mbox{}$ (3)
where $r_{t}(\theta)=\frac{\pi_{t}(a\mid c)}{\pi_{t-1}(a\mid c)}$ is the importance sampling ratio between the current policy $\pi_{t}(a\mid c)$ and the previous reference policy $\pi_{t-1}(a\mid c)$, $\hat{A}_{t}$ is the advantage function (sutton2018reinforcement), and the clip operator restricts the importance sampling ratio in the range $1-\epsilon,1+\epsilon]$.
2.3 RL for Text-to-Image Diffusion Models
The diffusion process can be viewed as an MDP $\left(\mathcal{S},\mathcal{A}\right$., $\left.\mathcal{P},\mathcal{R},\rho_{0}\right)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition kernel, $\mathcal{R}$ is the reward function, and $\rho_{0}$ is the distribution of initial state $\mathbf{s_{0}}$. In the context of text-to-image diffusion models, the MDP is defined as:
$\begin{split}\mathbf{s_{t}}=(\mathbf{c},t,\mathbf{x_{t}}),\;\;\pi_{\theta}(\mathbf{a_{t}}\mid\mathbf{s_{t}})=p_{\theta}(\mathbf{x_{t-1}}\mid\mathbf{x_{t}},\mathbf{c}),\;\;\mathcal{P}(\mathbf{s_{t+1}}\mid\mathbf{s_{t}},\mathbf{a_{t}})= \delta\big(\mathbf{c},\mathbf{a_{t}}\big),\;\;\mathbf{a_{t}}=\mathbf{x_{t-1}},\\ \rho_{0}(\mathbf{s_{0}})=\big(p(\mathbf{c}),\delta_{T},\mathcal{N}(0,\mathbf{I})\big),\;\;\mathcal{R}(\mathbf{s_{t}},\mathbf{a_{t}})= \begin{cases}r(\mathbf{x_{0}},\mathbf{c})&\text{if }t=0,\\ 0&\text{otherwise.}\end{cases}\end{split}$ (4)
The input state $\mathbf{s_{t}}$ is defined in terms of the context $\mathbf{c}$ (prompt features), and the sampled image at the given time-step $t$: $\mathbf{x_{t}}$. The policy $\pi_{\theta}$ is the diffusion model itself. The state transition kernel is a dirac delta function $\delta$ with the current sampled action $\mathbf{x}_{t}$ as the input. The reward is assigned only at the last step in the reverse diffusion process, when the final image is generated. The initial state $\rho_{0}$ corresponds to the last state in the forward diffusion process: $\mathbf{x}_{T}$.
2.4 PPO for Diffusion Fine-tuning
The objective function of RL fine-tuning for a diffusion policy $\pi_{\theta}$ can be defined as follows:
$J_{\theta}(\pi)=\mathbb{E}_{\tau\sim p(\tau\mid\pi_{\theta})}\left[\sum_{t=0}^{T}\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})\right]=\mathbb{E}_{\tau\sim p(\tau\mid\pi_{\theta})}\left[r(\mathbf{x}_{0},\mathbf{c})\right],$ (5)
where the trajectory $\mathbf{\tau}=\{\mathbf{x}_{T},\mathbf{x}_{T-1},\ldots,\mathbf{x}_{0}\}$ refers to the reverse diffusion process (Eq. 1), and the total reward of the trajectory is the reward of the final generated image $\mathbf{x}_{0}$ (Eq. 4). We ignore the KL-regularized version of the equation, which is commonly applied in the RLHF for LLM literature (zhong2024dpo; zeng2024token; rafailov2024direct), and proposed by fan2024reinforcement in the context of RL for diffusion models. As shown by black2023training, adding the KL-regularization term makes no empirical difference in terms of the final performance. The PPO objective is given as:
$J_{\theta}^{\mathrm{PPO}}(\pi)=\mathbb{E}\!\bigg[\!\sum_{t=0}^{T}\!\text{clip}\!\left(\!\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})},1\!-\!\epsilon,1\!+\!\epsilon\!\right)\!r(\mathbf{x}_{0},\mathbf{c})\!\bigg],$
where the clipping operation removes the incentive for the new policy $\pi_{\theta}$ to differ from the previous round policy $\pi_{\text{old}}$ (schulman2017proximal; black2023training).
3 REINFORCE vs. PPO: An Efficiency-Effectiveness Trade-Off
In this section, we explore the efficiency-effectiveness trade-off between two prominent reinforcement learning methods for diffusion fine-tuning: REINFORCE and PPO. Understanding this trade-off is crucial for selecting the appropriate algorithm given constraints on computational resources and desired performance outcomes.
In the context of text-to-image diffusion models, we aim to optimize the policy $\pi$ to maximize the expected reward $\mathcal{R}(x_{0:T},c)=r(x_{0},c)$. Our objective function is defined as:
$J_{\theta}(\pi)=\mathbb{E}_{c\sim p(C),x_{0:T}\sim p_{\theta}(x_{0:T}\mid c)}\left[r(x_{0},c)\right].$ (6)
REINFORCE for gradient calculation. For optimizing this objective, the REINFORCE policy gradient (also known as score function (SF)) (williams1992simple) provides the following gradient estimate:
$\begin{split}&\nabla_{\theta}J_{\theta}^{\mathrm{SF}}(\pi)\\ &=\mathbb{E}_{\mathbf{x}_{0:T}}\left[\nabla_{\theta}\log\left(\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{c}\right)\right)r\left(\mathbf{x}_{0},\mathbf{c}\right)\right]\\ &{}=\mathbb{E}_{\mathbf{x}_{0:T}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{c}\right)r\left(\mathbf{x}_{0},\mathbf{c}\right)\right],\end{split}$ (7)
where the second step follows from the reverse diffusion policy decomposition (Eq. 1).
In practice, a batch of trajectories is sampled from the reverse diffusion distribution, i.e., $\mathbf{x}_{0:T}\sim p_{\theta}(\mathbf{x}_{0:T})$, and a Monte-Carlo estimate of the REINFORCE policy gradient (Eq. 7) is calculated for the model update.
REINFORCE with baseline correction. To reduce variance of the REINFORCE estimator, a common trick is to subtract a constant baseline correction term from the reward function (greensmith2004variance; mohamed2020monte):
$\nabla_{\theta}J_{\theta}^{\mathrm{SFB}}(\pi)=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{c}\right)\left(r\left(\mathbf{x}_{0},\mathbf{c}\right)-b_{t}\right)\right].$ (8)
REINFORCE Leave-one-out (RLOO). To further reduce the variance of the REINFORCE estimator, RLOO samples $K$ diffusion trajectories per prompt ($\{\mathbf{x}^{k}_{0:T}\}\sim\pi(.\mid\mathbf{c}))$, for a better Monte-Carlo estimate of the expectation (kool2019buy; ahmadian2024back). The RLOO estimator is:
$\nabla_{\theta}J_{\theta}^{\mathrm{RLOO}}(\pi)=\mathbb{E}\left[\!K^{-1}\sum_{k=0}^{K}\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}^{k}_{t-1}\mid\mathbf{x}^{k}_{t},\mathbf{c}\right)\left(r\left(\mathbf{x}^{k}_{0},\mathbf{c}\right)-b_{t}\right)\!\right]\!.\\$ (9)
However, REINFORCE-based estimators have a significant disadvantage: they do not allow sample reuse (i.e., reusing trajectories collected from previous policies) due to a distribution shift between policy gradient updates during training. Sampled trajectories can only be used once, prohibiting mini-batch updates. This makes it sample inefficient.
To allow for sample reuse, the importance sampling (IS) trick can be applied (schulman2015trust; mcbook):
$J_{\theta}^{\textrm{IS}}(\pi)=\mathbb{E}_{c_{t}\sim p(C),a_{t}\sim\pi_{\text{old}}(a_{t}\mid c_{t})}\left[\frac{\pi_{\theta}(a_{t}\mid c_{t})}{\pi_{\text{old}}(a_{t}\mid c_{t})}\mathcal{R}_{t}\right],$ (10)
where $\pi_{\theta}$ is the current policy to be optimized, and $\pi_{\text{old}}$ is the policy from the previous update round. With the IS trick, we can sample trajectories from the current policy in a batch, store it in a temporary buffer, and re-use them to apply mini-batch optimization (schulman2017proximal).
Motivation for PPO. With the IS trick, the samples from the old policy can be used to estimate the policy gradient under the current policy $\pi_{\theta}$ (Eq. 7) in a statistically unbiased fashion (mcbook), i.e., in expectation the IS and REINFORCE gradients are equivalent (Eq. 10, Eq. 7). Thus, potentially, we can improve the sample efficiency of REINFORCE gradient estimation with IS.
While unbiased, the IS estimator can exhibit high variance (mcbook). This high variance may lead to unstable training dynamics. Additionally, significant divergence between the current policy $\pi_{\theta}$ and the previous policy $\pi_{\textrm{old}}$ can result in the updated diffusion policy performing worse than the previous one (schulman2015trust; achiam2017constrained). Next, we will prove this formally. We note that this result has previously been established by (achiam2017constrained) for the more general RL setting. In this work, we extend this finding to the context of diffusion model fine-tuning.
A key component of the proof relies on the distribution of states under the current policy, i.e., $d^{\pi}(s)$. In the case of diffusion models, the state transition kernel $P(s_{t+1}\mid s_{t},a_{t})$ is deterministic, because the next state consists of the action sampled from the previous state (Eq. 4), i.e., $P(s_{t+1}\mid s_{t},a_{t})=1$. While the state transition kernel is deterministic, the distribution of states is stochastic, given that it depends on the action at time $t$, which is sampled from the policy (Eq. 4). We define the state distribution as:
Definition 1.
Given the distribution over contexts $\mathbf{c}\sim p(\mathbf{C})$, the (deterministic) distribution over time $t=\delta(t)$, and the diffusion policy $\pi$, the state distribution at time $t$ is:
$p(\mathbf{s}_{t}\mid\pi)=p(\mathbf{c})\delta(t)\!\!\int_{\mathbf{x}_{t+1}}\hskip-15.6491pt\pi(\mathbf{x}_{t}|\mathbf{x}_{t+1},\mathbf{c},t)\pi(\mathbf{x}_{t+1}|\mathbf{c},t)\,\mathbf{dx}_{t+1}.$
Subsequently, the normalized discounted state visitation distribution can be defined as:
$d^{\pi}(\mathbf{s})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p(\mathbf{s}_{t}=\mathbf{s}\mid\pi).$ (11)
The advantage function is defined as: $A^{\pi_{k}}(\mathbf{s},\mathbf{a})=Q^{\pi_{k}}(\mathbf{s},\mathbf{a})-V^{\pi_{k}}(\mathbf{s})$ (sutton2018reinforcement). Given this, the monotonic policy improvement bound can be derived:
Theorem 3.1.
(achiam2017constrained) Consider a current policy $\pi_{k}$. Let $C^{\pi,\pi_{k}}=\max_{s\in S}|\mathbb{E}_{a\sim\pi(\cdot\mid s)}\left[A^{\pi_{k}}(s,a)\right]|$, and $\mathrm{TV}(\pi(\cdot\mid s),\pi_{k}(\cdot\mid s))$ represent the total variation distance between the policies $\pi(\cdot\mid s)$ and $\pi_{k}(\cdot\mid s)$, and $s$ be the current state. For any future policy $\pi$, we have:
$J(\pi)-J(\pi_{k})\geq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d^{\pi_{k}}}\left[\frac{\pi(a\mid s)}{\pi_{k}(a\mid s)}A^{\pi_{k}}(s,a)\right]-\frac{2\gamma C^{\pi,\pi_{k}}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d^{\pi_{k}}}\left[\mathrm{TV}(\pi(\cdot\mid s),\pi_{k}(\cdot\mid s))\right].$
A direct consequence of this theorem is that when optimizing a policy with the IS objective (Eq. 10), to guarantee that the new policy will improve upon the previous policy, the policies should not diverge too much. Therefore, we need to apply a constraint on the current policy. This can be achieved by applying the clipping operator in the PPO objective (Eq. 3(queeney2021generalized; achiam2017constrained; schulman2017proximal; gupta2024practical; gupta2024unbiased).
This gives rise to an efficiency-effectiveness trade-off between REINFORCE and PPO. REINFORCE offers greater computational and implementation efficiency due to its simplicity, but it comes at the cost of lower sample efficiency and potential suboptimal performance. In contrast, PPO is more computationally demanding and involves more complex hyper-parameter tuning, yet it achieves higher performance and reliable policy improvements during training.
We note that a similar trade-off analysis was performed in the context of RL fine-tuning for large language models (LLM) (ahmadian2024back). However, their analysis was limited to an empirical study, whereas we present a theoretical analysis in addition to the empirical analysis. To the best of our knowledge, we are the first to conduct such a study for diffusion methods.
4 Method: Leave-One-Out PPO (LOOP) for Diffusion Fine-tuning
We demonstrated the importance of PPO in enhancing sample efficiency and achieving stable improvements during training for diffusion fine-tuning. Additionally, we showcased the RLOO method’s effectiveness in reducing the variance of the REINFORCE method. In this section, we introduce our proposed method, LOOP, a novel RL for diffusion fine-tuning method. We start with highlighting the potential high-variance in the PPO objective.
The expectation in the PPO loss (Eq. 3) is typically estimated by sampling a single trajectory from the policy in the previous iteration $\pi_{old}$: $\mathbf{x}_{0:T}\sim\pi_{old}$ for a given prompt $c$:
$\mbox{}\hskip-2.84526pt\sum_{t=0}^{T}\text{clip}\left(\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})},1-\epsilon,1+\epsilon\!\right)r(\mathbf{x}_{0},\mathbf{c}).\!$ (12)
Even though the single sample estimate is an unbiased Monte-Carlo approximation of the expectation, it has high-variance (mcbook). Additionally, the IS term ($\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}$) can also contribute to high-variance of the PPO objective (swaminathan2015self; xie2023dropout). Both factors combined, can lead to high-variance, and unstable training of the PPO.
Taking inspiration from RLOO (Eq. 9), we sample $K$ independent trajectories from the previous policy for a given prompt $c$, and apply a baseline correction term from each trajectory’s reward, to reduce the variance of the estimator:
$\hat{J}_{\theta}^{\mathrm{LOOP}}(\pi)=\frac{1}{K}\!\sum_{i=1}^{K}\!\bigg[\!\sum_{t=0}^{T}\text{clip}\!\left(\frac{\pi_{\theta}(\mathbf{x}^{i}_{t-1}|\mathbf{x}^{i}_{t},c)}{\pi_{\text{old}}(\mathbf{x}^{i}_{t-1}|\mathbf{x}^{i}_{t},c)},1-\epsilon,1+\epsilon\right)\cdot\left(r(\mathbf{x}^{i}_{0},\mathbf{c})-b^{i}\right)\Big],$ (13)
where $\mathbf{x}^{i}_{0:T}\sim\pi_{old},\forall i\in[1,K]$. The baseline correction term $b^{i}$ reduces the variance of the gradient estimate, while being unbiased in expectation (gupta2024optimal; mohamed2020monte). A simple choice of baseline correction can be the average reward across the $K$ trajectories. However, it results in a biased estimator (kool2019buy). Therefore, we choose the leave-one-out average baseline, with average taken across all samples in the trajectory, except the current sample $i$, i.e.:
$b^{i}=\frac{1}{k-1}\sum_{j\neq i}r(\mathbf{x}^{j}_{0}).$ (14)
Originally RLOO sampling and baseline corrections were proposed in the context of REINFORCE, with a focus on on-policy optimization (ahmadian2024back; kool2019buy), whereas we are applying these in the off-policy step of PPO. We call this method leave-one-out PPO (LOOP).
Our approach is conceptually similar to the recently popular GRPO method for RL fine-tuning of LLMs (shao2024deepseekmath). Although our work was developed independently before GRPO gained widespread recognition, we do not include a head-to-head comparison.
Technically, the distinction lies in following aspects: (i) unlike GRPO, our formulation does not apply standard-deviation normalization in the denominator, as this has been shown to potentially harm performance in recent LLM fine-tuning via RL studies (liu2025understanding), (ii) similar to GRPO, we omit a KL penalty term, since our empirical experiments showed that it has little practical benefit. Furthermore, a recent study showed that on-policy RL implicitly constrains the updated policy to remain close to the base policy under a KL divergence measure, even without an explicit KL penalty term (shenfeld2025rl), and (iii) we ignore the generation-length normalization term. In the diffusion setting, this simplification is further justified by the fact that the sequence length of the reverse diffusion process is fixed across generations, rendering length normalization unnecessary.
5 Experimental Setup
Benchmark. Text-to-image diffusion and language models often fail to satisfy an essential reasoning skill of attribute binding. Attribute binding reasoning capability refers to the ability of a model to generate images with attributes such as color, shape, texture, spatial alignment, (and others) specified in the input prompt. In other words, generated images often fail to bind certain attributes specified in the instruction prompt (huang2023t2i; ramesh2022hierarchical; fu2024enhancing). Since attribute binding seems to be a basic requirement for useful real-world applications, we choose the T2I-CompBench benchmark (huang2023t2i), which contains multiple attribute binding/image compositionality tasks, and its corresponding reward metric to benchmark text-to-image generative models. We also select two common tasks from prior RL for diffusion work: improving aesthetic quality of generation, and image-text semantic alignment (black2023training; fan2024reinforcement). To summarize, we choose the following tasks for the RL optimization: (i) Color, (ii) Shape, (iii) Texture, (iv) 2D Spatial, (v) Numeracy, (vi) Aesthetic, (vii) Image-text Alignment. For all tasks, the prompts are split into training/validation prompts. We report the average reward on both training and validation split.
Model. As the base diffusion model, we use Stable diffusion V2 (rombach2022high), which is a latent diffusion model. For optimization, we fully update the UNet model, with a learning rate of $1e^{-5}$. We also tried LORA fine-tuning (hu2021lora), but the results were not satisfactory, so we update the entire model instead.
6 Hyperparameter and Implementation Details
For REINFORCE (including REINFORCE with baseline correction term), PPO, and LOOP the number of denoising steps ($T$) is set to 50. The diffusion guidance weight is set to 5.0. For optimization, we use AdamW Loshchilov2017DecoupledWD with a learning rate of $1e^{-5}$, and the weight decay of 1e41e{-4}, with other parameters kept at the default value. We clip the gradient norm to $1.0$. We train all models using 8 A100 GPUs with a batch size of 4 per GPU. The clipping parameter $\epsilon$ for PPO, and LOOP is set to $1e^{-4}$.
7 Results and Discussion
7.1 REINFORCE vs. PPO efficiency-effectiveness trade- off
We present our empirical results on the efficiency-effectiveness trade-off between REINFORCE and PPO. Our evaluation compares the following methods: the REINFORCE policy gradient for diffusion fine-tuning (Eq. 7); the REINFORCE policy gradient with a baseline correction term (REINFORCE w/ BC), detailed in Eq. 8, where the baseline term is the average reward for the given prompt (black2023training), and the PPO objective for diffusion fine-tuning, which incorporates importance sampling and clipping, as outlined in Eq. 3. This PPO objective is equivalent to the DDPO objective in the original RL for diffusion method (black2023training).
Figure 2 shows the training reward over epochs for the attributes: Color, Shape, and Texture from the T2I-CompBench benchmark, and training reward from optimizing the aesthetic model. Results are averaged over 3 runs. It is clear that REINFORCE policy gradient is not effective in terms of performance, as compared to other variants. Adding a baseline correction term indeed improves the training performance, validating the effectiveness of baseline in terms of training performance, possibly because of reduced variance. PPO achieves the highest training reward, validating the effectiveness of importance sampling and clipping for diffusion fine-tuning.
   Color    Shape    Texture    Aesthetic
Reward
Refer to caption Refer to caption Refer to caption Refer to caption
    Epochs     Epochs     Epochs     Epochs
Figure 2: Evaluating REINFORCE vs. PPO trade-off by comparing: REINFORCE (Eq. 7), REINFORCE with baseline correction term (Eq. 8), and PPO (Eq. 3). We evaluate on the T2I-CompBench benchmark over three image attributes: Color, Shape, and Texture. We also compare on the aesthetic task. Y-axis corresponds to the training reward, x-axis corresponds to the training epoch. Results are averaged over 3 runs; shaded areas indicate 80% prediction intervals.
We also evaluate the performance on a separate validation set. For each validation prompt, we generate 10 independent images from the diffusion policy, and average the reward, finally averaging over all evaluation prompts. The validation results are reported in Table 1. The results are consistent with the pattern observed with the training rewards, i.e., REINFORCE with baseline provides a better performance than plain REINFORCE, suggesting that baseline correction indeed helps with the final performance. Nevertheless, PPO (DDPO) still performs better than REINFORCE.
Table 1: Comparing REINFORCE with DDPO on the T2I-CompBench benchmark over three image attributes: Color, Shape, and Texture. We report average reward on unseen test set (higher is better). For each prompt, average rewards over 10 independent generated images are calculated.
Method Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
REINFORCE 0.6438 0.5330 0.6359
REINFORCE w/ BC 0.6351 0.5347 0.6656
DDPO 0.6821 0.5655 0.6909
We now have empirical evidence supporting the efficiency-effectiveness trade-off discussed in Section 3. From these results, we can conclude that fine-tuning text-to-image diffusion models is more effective with IS and clipping from PPO, or baseline corrections from REINFORCE. This bolsters our motivation for proposing LOOP as an approach to effectively combine these methods.
7.2 Evaluating LOOP
Next we discuss the results from our proposed RL for diffusion fine-tuning method, LOOP.
Performance during training. Figure 3 shows the training reward curves for different tasks, against number of epochs. LOOP outperforms DDPO (black2023training) across all seven tasks consistently throughout training. This establishes the effectiveness of sampling multiple diffusion trajectories per input prompt, and the leave-one-out baseline correction term (Eq. 9) during training. Training reward curve is smoother for the aesthetic task, as compared to tasks from the T2I-CompBench benchmark. We hypothesise that improving the attribute binding property of diffusion model is a harder task than improving the aesthetic quality of generated images.
Color Shape Texture
Reward
Refer to caption Refer to caption Refer to caption
    Spatial     Image-Text Alignment     Aesthetic
Reward
Refer to caption Refer to caption Refer to caption
Figure 3: Comparing DDPO (referenced as PPO) with the proposed LOOP on the T2I-CompBench benchmark with respect to image attributes: Color, Shape, Texture, and Spatial relationship. We also report results on aesthetic preference and image–text alignment tasks (black2023training). The y-axis shows training reward, and the x-axis shows training epoch. Results are averaged over three independent runs; shaded areas denote 80% prediction intervals.
Table 2: Comparing the performance of the proposed LOOP method with state-of-the-art baselines on the T2I-CompBench benchmark over image attributes such as Color, Shape, Texture, Spatial relation, and Numeracy. The metrics in this table are average reward on an unseen test set (higher is better). For each prompt we generate and average rewards across 10 different generated images.
Model Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ Spatial $\uparrow$ Numeracy $\uparrow$
Stable v1.4 (rombach2022high) 0.3765 0.3576 0.4156 0.1246 0.4461
Stable v2 (rombach2022high) 0.5065 0.4221 0.4922 0.1342 0.4579
Composable v2 (liu2022compositional) 0.4063 0.3299 0.3645 0.0800 0.4261
Structured v2 (feng2022training) 0.4990 0.4218 0.4900 0.1386 0.4550
Attn-Exct v2 (chefer2023attend) 0.6400 0.4517 0.5963 0.1455 0.4767
GORS unbiased (huang2023t2i) 0.6414 0.4546 0.6025 0.1725
GORS (huang2023t2i) 0.6603 0.4785 0.6287 0.1815 0.4841
DDPO (black2023training) 0.6821 0.5655 0.6909 0.1961 0.5102
LOOP ($k=3$) 0.7515 0.6220 0.7353 0.1966 0.5242
LOOP ($k=4$) 0.7859 0.6676 0.7518 0.2136 0.5422
Table 2 reports the average rewards on the test set across various tasks from the T2I-CompBench benchmark. For each prompt, we generate 10 different images and calculate the average rewards. LOOP consistently outperforms DDPO (black2023training) and other strong supervised learning-based baselines across all tasks. Notably, LOOP achieves relative improvements of 18.1% and 15.2% over DDPO on shape and color attributes, respectively.
For the aesthetic and image-text alignment objectives, the validation rewards are reported in Table 3. LOOP results in a 15.4% relative improvement over PPO for the aesthetic task, and a 2.4% improvement over PPO for the image-text alignment task.
Impact of number of independent trajectories ($k$). The LOOP variant with number of independent trajectories $K=4$ performs the best across all tasks, followed by the variant $K=3$. This is intuitive given that Monte-Carlo estimates get better with more number of samples (mcbook). Surprisingly, the performance of the variant with $K=2$ is comparable to PPO.
8 Qualitative Examples
For a qualitative evaluation of the attribute-binding reasoning ability, we present some example image generations from SD, DDPO, and LOOP in Figures 1, 4, and 5.
In Figure 1 qualitative examples of the attribute binding task are presented. In the example in the first column of Figure 1, input prompt specifies a black ball with a white cat. Stable diffusion (SD) and PPO fail to bind the color black with the generated ball, whereas LOOP successfully binds that attribute. Similarly, in the third column, SD and PPO fail to bind the hexagon shape attribute to the watermelon, whereas LOOP manages to do that. In the fourth column, SD and PPO fail to add the horse object itself, whereas LOOP adds the horse with the specified black color, and flowing cyan patterns.
Figure 4 highlights improvements in aesthetic quality of the generated images. Compared to SD v2 and PPO, LOOP produces sharper, more coherent compositions with balanced lighting and color tone. For example, in the second column (“a cat”) and in the fourth column (“butterfly”), LOOP enhances realism and contrast while preserving overall artistic intent.
Finally, Figure 5 presents additional qualitative examples that emphasize both binding and aesthetics. LOOP accurately binds challenging color-object pairs (e.g., teal branch, pink cornfield) while producing more visually appealing and natural results. PPO and SD v2 often miss attribute alignment or produce dull, less cohesive scenes.
Table 3: Comparing the performance of LOOP with DDPO on the aesthetic and image-text alignment tasks. Higher values are better.
Method Aesthetic $\uparrow$ Image Align. $\uparrow$
DDPO (black2023training) 6.8135 20.466
LOOP ($k=2$) 6.8617 20.788
LOOP ($k=3$) 7.0772 20.619
LOOP ($k=4$) 7.8606 20.909
SD v2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PPO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LOOP Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A puppy dog”
“A cat”
“Butterfly”
“Bright yellow sunflower in a green field”
“Crystal clear mountain lage reflecting snow-capped peaks”
Figure 4: LOOP improves aesthetic quality. Qualitative examples are presented from images generated via: Stable Diffusion 2.0 (first row), PPO (second row), and LOOP $k=4$ (third row). LOOP consistently generates more aesthetic images, as compared to PPO and SD.
SD v2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PPO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LOOP Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A white feather on a black velvet surface” “A neon orange owl sitting on a teal branch” “Pink bridge over a glowing blue river” “A pink cornfield”
“A yellow flower in a green vase”
Figure 5: Additional qualitative examples presented from images generated via Stable Diffusion 2.0 (first row), PPO (second row), and LOOP $k=4$ (third row). LOOP consistently generates more aesthetic images, as compared to PPO and SD (first, third, and fifth prompt). LOOP also binds the color attribute (teal branch in second example, and pink cornfield in the forth example), where SD and PPO fail.
9 Conclusion
We have studied the efficiency-effectiveness trade-off between two fundamental RL methods for diffusion fine-tuning: REINFORCE and PPO. Our analysis, both theoretical and empirical, demonstrates that while REINFORCE is computationally efficient and easier to implement, it suffers from high variance and sample inefficiency compared to PPO. PPO, though more effective, comes with significant computational overhead, requiring three models in memory simultaneously and involving sensitive hyperparameter tuning.
Building on these insights, we have introduced LOOP, a novel RL method for diffusion fine-tuning that combines variance reduction techniques from REINFORCE (multiple trajectory sampling and leave-one-out baseline correction) with the robustness and sample efficiency of PPO (importance sampling and clipping). Our empirical evaluation on the T2I-CompBench benchmark demonstrates that LOOP achieves substantial improvements over both the base Stable Diffusion model and the state-of-the-art PPO method across multiple tasks, including attribute binding (color, shape, texture, spatial relationships), aesthetic quality, and image-text alignment.
Quantitatively, LOOP (k=4) achieves substantial improvements over PPO across all evaluated tasks. On the T2I-CompBench benchmark, LOOP achieves relative improvements of $\mathbf{18.1}\%$ on shape binding, $\mathbf{15.2}\%$ on color binding, $\mathbf{8.8}\%$ on texture binding, and $\mathbf{8.9}\%$ on spatial reasoning. LOOP also improves aesthetic quality by $\mathbf{15.4}\%$ and image-text alignment by $\mathbf{2.2}\%$. Qualitatively, as shown in Figures 1, 4, and 5, LOOP successfully binds attributes that previous methods fail to capture, while also producing more visually coherent and aesthetic images.
A limitation of LOOP is the increased computational cost from sampling multiple diffusion trajectories per prompt, which leads to longer training times compared to standard PPO. Future work could explore adaptive sampling strategies to reduce this overhead while maintaining LOOP’s effectiveness, extend the method to other diffusion architectures and modalities, or investigate the integration of human preference modeling for better alignment with real-world objectives.